# Keeping Up with the Models: Online Deployment and Routing of LLMs at Scale

### v1

@@ Abstract
The rapid pace at which new large language models (LLMs) appear—and older ones become obsolete—forces LLM service providers to juggle a streaming inventory of models while respecting tight deployment capacity and per-query cost budgets. We cast the reality as an online decision problem that couples stage-wise deployment, made at fixed maintenance windows, with per-query routing among the models kept live. We introduce StageRoute, a hierarchical algorithm that (i) optimistically selects up to $M_{\max}$ models for the next stage using reward upper-confidence and cost lower-confidence bounds, then (ii) solves a budget-constrained bandit sub-problem to route each incoming query. We prove that StageRoute achieves a regret of $\tilde{\mathcal{O}}(T^{2/3})$ and provide a matching lower bound, thereby establishing its near-optimality. Moreover, our experiments confirm the theory, demonstrating that StageRoute performs close to the optimum in practical settings.
@@ 1 Introduction
The proliferation of large language models (LLMs) has revolutionized a broad array of applications, delivering unprecedented advances in natural-language understanding and generation [35, 11, 44, 1, 14, 40]. Yet LLMs differ greatly in both performance and cost: some offer state-of-the-art capabilities at a premium, while others are more affordable but less effective. Practitioners must therefore strike a careful balance between accuracy and expenditure when deploying models. This need has given rise to the problem of LLM routing [18, 24], in which a system decides—query by query—which model to invoke, aiming to maximize task quality subject to real-time cost constraints.
The LLM ecosystem is evolving at breakneck speed: new models appear continuously, each with its own accuracy–latency profile and cost structure (for example, per-token API pricing) [21]. The challenge is even sharper in production settings, where the pool of available models changes over time and operators must honor additional constraints such as throughput limits and deployment quotas. Azure OpenAI Service, for instance, caps each resource at 32 standard deployments and 5 fine-tuned deployments by default, and enforces model-specific rate ceilings—GPT-4.1 is limited to 1,000 requests per minute (RPM) and 1 million tokens per minute (TPM) [32].
Motivated by these challenges, we introduce an online decision framework that tackles two intertwined tasks (see Figure 1 for an overview): (1) Stage-wise deployment: At fixed update intervals, decide which LLMs to add or retire, accounting for evolving prices and non-trivial initialization or licensing fees; and (2) Per-query routing: Between updates, assign each user query to a currently deployed model to maximize quality while honoring strict cost budgets. Unlike approaches that assume a static model pool or rely on fully offline retraining, our framework explicitly allows new LLMs to arrive over time. Within each update window it can (i) invest the one-time cost to deploy a promising newcomer discovered during exploration, or (ii) drop a persistently underperforming or overpriced model. This adaptability lets service providers keep pace with rapid LLM advances, maintain service-level quality, and stay within tight resource budgets.
This setting is challenging on several fronts. First, because new—and potentially superior—models continuously appear, any algorithm must adapt its strategy as the candidate pool evolves. Second, updates to the deployed model set occur only at fixed intervals, creating a hierarchical decision structure: long-horizon deployment choices determine the tactical, per-query routing possibilities between updates. Finally, the problem combines a global cost budget, per-model capacity limits, and the combinatorial task of selecting a small subset of models under uncertainty—dramatically increasing its complexity. Standard online-learning tools such as multi-armed bandits or online convex optimization do not natively handle this mix of dynamic availability, staged deployment, and coupled budget-and-capacity constraints.
To tackle these challenges, we introduce StageRoute, an algorithm that mirrors the problem’s hierarchical structure (see Figure 1). At each update boundary, StageRoute enters a model deployment phase, selecting the set of models for the next stage. Using optimistic upper-confidence bounds (UCB) for rewards and conservative lower-confidence bounds (LCB) for costs, it picks the subset that maximizes expected utility while respecting the global budget, per-model capacity limits, and the $M_{\max}$ deployment cap. During the ensuing stage, every incoming query triggers StageRoute’s request routing phase: a linear program over the currently deployed models produces a routing distribution that maximizes estimated quality under the same cost and capacity constraints, and the query is dispatched by sampling from this distribution. This two-level design links strategic stage-level deployment to fine-grained, adaptive routing, allowing the system to integrate new information both across and within stages.
In summary, our contributions in this paper are threefold:
$\bullet$ Problem formulation and algorithm. We formalize the online LLM routing problem with dynamic arrivals and staged deployment and present StageRoute, a hierarchical algorithm that decouples strategic deployment from per-query routing via confidence bounds.
$\bullet$ Theoretical guarantees. We upper bound StageRoute’s regret against an offline oracle and prove a matching $\Omega(T^{2/3})$ lower bound for any online algorithm.
$\bullet$ Empirical evaluation. Experiments show that StageRoute matches a strong oracle within tight budgets and stays robust across key parameters settings.
@@
Figure 1: Overview of the StageRoute workflow. Newly released LLMs (green) continuously enter the candidate pool. At each scheduled update point, StageRoute selects up to $M_{\max}$ to deploy (blue). Between updates, every incoming query is routed among the currently deployed models (orange). This two-level loop lets StageRoute assimilate fresh models, stay within cost constraints, and adapt routing decisions in real time.
2 Related Work
LLM Routing. The central aim of LLM routing is to strike the best balance between task performance (e.g., response quality or accuracy) and operational metrics such as cost and latency [18, 4]. Existing work follows three main architectural patterns. Ensemble strategies query multiple models in parallel to boost robustness, at the expense of higher cost and latency [42, 26]. Cascade strategies issue queries sequentially—typically starting with a cheaper model and escalating only when necessary—thereby reducing cost but sometimes increasing latency [12, 22, 45, 4].Direct-routing strategies train a policy or classifier that selects a single LLM per query [34, 21, 47, 50]. Benchmark suites such as RouterBench [24] facilitate systematic comparison of these approaches. Related work on mixture-of-experts (MoE) models explores routing within a single large model [19, 20, 36]. More recently, bandit formulations have been applied to static LLM routing [43, 17, 31, 33]. Most prior studies, however, assume a fixed set of available models and focus solely on per-query decisions. In contrast, we explicitly model dynamic LLM arrivals, enforce staged deployment updates—where the active model set can change only at periodic intervals—and respect tight constraints on budget, throughput, and the maximum number of simultaneously deployed models ($M_{\max}$).
Multi-armed Bandits. Our formulation builds on the multi-armed-bandit (MAB) paradigm, in which an agent sequentially selects among several actions (“arms”) with initially unknown rewards, learning to maximize its cumulative payoff through exploration and exploitation [7, 39]. Two MAB extensions are especially pertinent: (i) Bandits with knapsacks (BwK). Here each arm pull yields a reward and consumes limited resources from one or more budgets; the objective is to maximize total reward without overspending [10, 5, 25, 28]. Our long-term cost constraint fits naturally into this framework. (2) Streaming bandits. In this setting new arms arrive over time—often under memory or attention limits—so the agent must adapt to a continually expanding action set [6, 27, 3, 41, 30, 38, 49]. The steady appearance of new LLMs places our problem squarely in this category. Our scenario combines both challenges and introduces an additional wrinkle: staged deployment. The set of usable LLMs (arms) can change only at discrete update points, creating a hierarchical decision process in which infrequent, strategic deployment choices constrain frequent, tactical routing decisions. This structure is largely absent from existing BwK or streaming-bandit literature and motivates the two-level algorithm we develop.
@@ 3 System Model
We study an online LLM routing framework: at each round $t$, a query arrives and must be routed to a suitable large language model (LLM). New models may appear at any time, yet they can be activated only at discrete deployment intervals (Figure 1). We describe each component of the model in detail.
@@
Rounds (Queries). Let $T$ denote the total number of user queries, indexed by $[T]=\{1,2,\dots,T\}$. At each time step $t\in[T]$, a query arrives and is immediately rerouted to an LLM $m_{t}$ chosen from the currently deployed models according to the algorithm’s policy.
@@
LLM Pool and Deployment Schedule. Let $\mathcal{M}_{t}$ be the set of all LLM models that exist and could, in principle, be deployed by time $t$. Each model $m$ has an availability time $t_{m}$; hence $m\in\mathcal{M}_{t}$ exactly when $t\geq t_{m}.$ Deployment changes occur only at fixed intervals. Divide the horizon of $T$ queries into $K$ equal-length stages of $T/K$ rounds (assume $T$ is divisible by $K$). The start of stage $k$ is $\tau_{k}=(k-1)T/K+1,k=1,\dots,K.$ Let $M_{\max}$ be the maximum number of models that the algorithm can deploy simultaneously. At each update point $\tau_{k}$, the algorithm $\mathcal{A}$ selects a deployed set $\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$ and $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$, which then remains fixed for all $t\in[\tau_{k},\tau_{k+1}).$ Queries arriving during that stage must be routed to models in the active set $\mathcal{D}_{k}(\mathcal{A})$. Thus $\mathcal{M}_{t}$ is the available pool at time $t$, while $\mathcal{D}_{k}(\mathcal{A})$ is the active subset that can actually serve queries during stage $k$.
@@
Operational Performance and Cost. Each model $m$ has an unknown performance distribution $\nu_{m}(\cdot)$ supported on $[0,1]$. When $m$ is selected at time $t$, the observed score $r_{t}\in[0,1]$ is drawn from $\nu_{m}(\cdot)$ with mean $\mu_{m}\;=\;\mathbb{E}_{x\sim\nu_{m}(\cdot)}[x].$ Invoking model $m$ on a query also incurs an operational cost $c_{m_{t}}$ that combines:
(i) Input Cost: $c_{m_{t}}^{(\mathrm{in})}=\bigl{(}\text{\# tokens of input at time }t\bigr{)}\times p_{\mathrm{in}},$ where $p_{\mathrm{in}}$ is the per-token input price;
(ii) Output Cost: $c_{m_{t}}^{(\mathrm{out})}=\bigl{(}\text{\#tokens of response by }m\text{ for query }t\bigr{)}\times p_{\mathrm{out}},$ with output length drawn from a model-specific distribution $\xi_{m}(\cdot)$ and unit price $p_{\mathrm{out}}$.
@@
Thus the total cost is $c_{m_{t}}=c_{m_{t}}^{(\mathrm{in})}+c_{m_{t}}^{(\mathrm{out})}.$ Because the output token count—and therefore $c_{m_{t}}^{(\mathrm{out})}$—depends on the specific model and query and is sampled from $\xi_{m}(\cdot)$, the total operational cost $c_{m_{t}}$ for a query handled by model $m_{t}$ is itself a random variable. Operational cost $c_{m_{t}}$ is inherently bounded by per-token pricing and practical limits on sequence length and generation (e.g., context-window and token caps). Hence we assume $c_{m_{t}}\in[c_{1},c_{2}]$ for known constants $0<c_{1}\leq c_{2}<\infty$.
@@
Cost Constraints and Capacity Limits. When a query arrival at time $t$, with stage index $k$ such that $\tau_{k}\leq t<\tau_{k+1}$, the algorithm $\mathcal{A}$ selects a model $m_{t}\in\mathcal{D}_{k}(\mathcal{A})$. It then observes the reward $r_{t}\sim\nu_{m_{t}}(\cdot)$ and incurs cost $c_{m_{t}}$. The goal is to maximize average reward subject to (1) Long-term budget: $\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\bigl{[}\sum_{t=1}^{T}c_{m_{t}}\bigr{]}\leq b.$ (2) Per-model capacity: For each deployed LLM model $m$, we specify a load cap $\alpha_{m}$. Let $p_{t}(m)$ be the probability that the routing policy assigns the query at time $t$ to model $m$. While $t$ lies in stage $k$, $p_{t}(m)\leq\alpha_{m},\forall m\in\mathcal{D}_{k}(\mathcal{A}).$
@@
Here, $p_{t}(m)$ is the probability that the routing policy assigns the query at time $t$ to model $m$. This constraint caps the instantaneous load share each model may receive. For a deterministic decision $m_{t}$, $p_{t}(m_{t})=1$ and $p_{t}(m)=0$ for $m\neq m_{t}$; the requirement then reduces to $\alpha_{m_{t}}\geq 1$. Such limits reflect real-world restrictions—bandwidth, licensing, or throughput—preventing any single model from handling an excessive fraction of queries within a stage. Hence the deployed set’s aggregate capacity must be large enough to serve every arrival, a condition we formalize next.
@@
The constraints are indeed feasible. At every time $t$, one can select a subset $\mathcal{S}\subseteq\mathcal{M}_{t}$ with $|\mathcal{S}|\leq M_{\max}$ such that $\sum_{m\in\mathcal{S}}\alpha_{m}\geq 1$. The budget $b$ is large enough to admit non-trivial routing. When required, we assume Slater’s condition holds, guaranteeing strong duality for the associated optimization problems.
Performance Maximization and Regret. The goal is to maximize the expected cumulative performance, $\mathbb{E}[\sum_{t=1}^{T}\mu_{m_{t}}]$ subject to:
$\triangleright$ Deployment Choice: At each update $\tau_{k}$, select a deployed set $\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$ with $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$.
$\triangleright$ Model Selection: For $t\in[\tau_{k},\tau_{k+1})$, choose $m_{t}\in\mathcal{D}_{k}(\mathcal{A})$ using probabilities $p_{t}(m)$ that sum to $1$.
$\triangleright$ Capacity Constraint: Ensure $p_{t}(m)\leq\alpha_{m}$ for every deployed model $m.$
$\triangleright$ Budget Constraint: Maintain $\lim_{T\to\infty}\frac{1}{T}\mathbb{E}[\sum_{t=1}^{T}c_{m_{t}}]\leq b$.
@@
[…]
@@
Here, $\Delta(\mathcal{S})$ is the set of probability distributions over $\mathcal{S}$; $p(m)$ is the probability of selecting model $m$; and $\mathbb{E}[c_{m}]$ is its expected cost. The support constraint $|\text{supp}(p)|\leq M_{\max}$ limits the number of models with positive probability to $M_{\max}$. If no feasible distribution exists—or if $(\mathcal{S}=\varnothing)$—we set $V(b,\mathcal{S})=0$.
@@
4 StageRoute: Stage-Based LLM Deployment and Routing Algorithm

### v2

# Near-Optimal Online Deployment and Routing for Streaming LLMs

@@ Abstract
The rapid pace at which new large language models (LLMs) appear, and older ones become obsolete, forces providers to manage a streaming inventory under a strict concurrency cap and per-query cost budgets. We cast this as an online decision problem that couples stage-wise deployment (at fixed maintenance windows) with per-query routing among live models. We introduce StageRoute, a hierarchical algorithm that (i) optimistically selects up to $M_{\max}$ models for the next stage using reward upper-confidence and cost lower-confidence bounds, and (ii) routes each incoming query by solving a budget- and throughput-constrained bandit subproblem over the deployed set. We prove a regret of $\tilde{\mathcal{O}}(T^{2/3})$ with a matching lower bound, establishing near-optimality, and validate the theory empirically: StageRoute tracks a strong oracle under tight budgets across diverse workloads.
@@ 1 Introduction
The proliferation of LLMs has transformed a broad array of applications, delivering unprecedented advances in natural-language understanding and generation (Radford et al., 2019; Brown et al., 2020; Wang et al., 2023b; (2023), 2023; Chowdhery et al., 2023; Touvron et al., 2023). Yet LLMs differ markedly in both performance and cost: some offer state-of-the-art capabilities at a premium, while others are more affordable but less effective. Practitioners therefore face a continual accuracy-expenditure tradeoff when deciding which models to operate and when to use them. This has motivated LLM routing (Ding et al., 2024; Hu et al., 2024), where a system chooses, query by query, which model to invoke to maximize task quality under cost constraints. However, focusing solely on per-query routing overlooks a more fundamental decision that precedes it: which models are deployed at all.
In practice, the operational landscape is unusually fluid. New models arrive continuously with distinct accuracy, latency, and pricing profile (Feng et al., 2025), while production systems must respect hard limits such as rate ceilings and deployment quotas. For example, Azure OpenAI Service caps each resource at 32 standard and 5 fine-tuned deployments by default, and enforces model-specific rate ceilings (e.g., for GPT-4.1: 1,000 requests per minute (RPM) and 1M tokens per minute (TPM)) (Microsoft Azure, 2025). This confluence of a dynamic model pool and strict operational caps recasts the problem into two timescales: a slower stage-wise deployment process that decides which models stay alive under a concurrency cap, and a faster per-query routing process that assigns each request among the currently deployed models while meeting budget and throughput constraints. The deployment choice is foundational, since it determines the entire action space for any routing policy. Table 1 maps recent LLM routing systems to three axes and highlights a gap in current approaches.
@@
Figure 1: The StageRoute workflow. Newly released LLMs (green) continually enter the candidate pool. At each scheduled update point, StageRoute deploys up to $M_{\max}$ models (blue). Between updates, each query is routed among the current deployment (orange). This two-level loop assimilates fresh models, enforces cost/throughput constraints, and adapts routing in real time.
We study this setting as an online decision problem that couples two intertwined choices (Figure 1): (1) stage-wise deployment at fixed update points, where the operator decides which models to deploy for the next stage subject to a hard concurrency cap and deployment costs. This high-stakes decision defines the action set for the subsequent execution of (2) per-query routing, where each incoming query is sent to one of the currently deployed models to maximize quality while obeying a long-term cost budget and per-model throughput limits. Unlike approaches that assume a static model pool or rely on fully offline retraining, our framework admits streaming arrivals of new LLMs and enforces active-set replacement: admitting a newcomer may require evicting an incumbent for the rest of the stage. This mirrors real service constraints while enabling continual adaptation.
Three features make this problem technically distinct. First, the hard concurrent-deployment cap induces an irreversible exploration-exploitation tradeoff: activating an uncertain model can mean dropping a known, reliable one for an entire stage. Second, decisions occur on two timescales: infrequent, strategic deployment choices constrain frequent, tactical per-query routing. Third, the system must jointly respect a long-term cost budget and per-model throughput limits while selecting a small operational subset under uncertainty. Classical multi-armed bandits (MAB), budgeted formulations (BwK), combinatorial bandits (CMAB), and standard streaming bandits each capture parts of this picture, but none natively address the combination of dynamic availability, staged commitment, and an explicit concurrency cap on the active set.
To address these challenges, we introduce StageRoute, an algorithm that mirrors the problem’s hierarchy (Figure 1). At each update boundary, a deployment phase selects the active set for the next stage using optimistic estimates of model quality (UCB) and conservative estimates of cost (LCB), honoring the global budget, per-model throughput limits, and the $M_{\max}$ concurrency cap. Within the stage, each query triggers a routing phase: a linear program over the currently deployment returns a distribution that maximizes estimated quality under the same constraints, and the query is dispatched by sampling accordingly. This two-level loop links strategic deployment to fine-grained, adaptive routing, allowing the system to assimilate new information both across and within stages. [^1] Relative to nearby bandit frameworks: static-pool routing assumes fixed arms; BwK models consumable budgets but not stage-level active-set replacement; CMAB selects superarms from a fixed base set without streaming arrivals or stage commitment; streaming bandits allow arrivals but do not couple stage-level support selection with per-query routing under both a budget and per-model capacity. See Appendix A for more details.
Our contributions in this paper are summarized as follows:
$\bullet$ Problem formulation. To our knowledge, we are the first to formalize the online LLM deployment and routing problem with streaming arrivals, explicitly modeling a hard concurrency cap, one-time deployment costs, per-model throughput limits, and a long-term cost budget, with stage-level commitment and per-query routing.
$\bullet$ Algorithm. We introduce StageRoute, which (i) selects an active set at each update using optimistic performance (UCB) and conservative cost (LCB) estimates under the budget, throughput limits, and the $M_{\max}$ concurrency cap, and (ii) routes each query by solving a budget–throughput LP over the currently deployed models. The design is modular: the routing step can incorporate contextual estimators when features are available, while the deployment step remains unchanged; throughput limits naturally throttle load to mitigate latency spikes.
$\bullet$ Theoretical guarantees. We prove a regret bound of $\widetilde{\mathcal{O}}\!\big(\sqrt{M_{\max}KT}\big)+\widetilde{\mathcal{O}}\!\big(NT/(M_{\max}K)\big)$, where $T,K,M_{\max},N$ are the numbers of queries, update stages, the concurrency cap and arriving models, respectively. The first term captures the statistical learning cost of routing within the deployed set; it grows with the number of active models $M_{\max}$, stages $K$, and horizon $T$. The second term is a structural model-discovery bottleneck that quantifies the difficulty of discovering strong newcomers when only $M_{\max}$ models can be live across $K$ stages as $N$ models arrive. Balancing the two yields a near-optimal $\widetilde{\mathcal{O}}(T^{2/3})$ order, and we give a matching $\Omega(T^{2/3})$ lower bound via a staged-arrival construction. Analytically, we use LP-duality sensitivity with an explicit support (active-set) constraint and a regret decomposition that separates routing and deployment across stages.
$\bullet$ Empirical evaluation. Simulations show StageRoute tracks a strong oracle under tight budgets and is robust across key parameters. We evaluate on true per-query scores and costs (RouterBench) across diverse queries, tasks, and languages, demonstrating effectiveness in realistic settings.
Table 1: Comparison of LLM routing frameworks. StageRoute is the first to address the full real-world setting: a dynamic model pool with streaming arrivals, paired with dynamic stage-wise deployment under a concurrency cap ($M_{\max}$) and cost- and throughput-aware routing.
Approach Streaming LLM Models Dynamic Deployment (with $M_{\max}$ cap) Cost & Budget Aware Throughput Limits Source
LLM-Blender Jiang et al. (2023)
AutoMix Aggarwal et al. (2024)
Hybrid-LLM Ding et al. (2024)
Zooter Lu et al. (2024)
RouterDC Chen et al. (2024)
TensorOpera Router Stripelis et al. (2024)
RouteLLM Ong et al. (2025)
MESS+ Woisetschläger et al. (2025)
UniRoute Jitkrittum et al. (2025)
CSCR Shirkavand et al. (2025)
StageRoute (ours) This paper
@@ 2 System Model
We study an online LLM routing framework: at each round $t$, a query arrives and must be routed to a suitable LLM. New models may appear at any time, yet they can be activated only at discrete deployment intervals (Figure 1). We describe each component of the model in detail.
@@
Rounds (Queries). Let $T$ denote the total number of user queries, indexed by $T]=\{1,2,\dots,T\}$. At each time step $t\in[T]$, a query arrives and is immediately routed to an LLM $m_{t}$ chosen from the currently deployed models according to the algorithm’s policy.
@@
LLM Pool and Deployment Schedule. Let $\mathcal{M}_{t}$ be the set of all LLM models that exist and could, in principle, be deployed by time $t$. Each model $m$ has an availability time $t_{m}$; hence $m\in\mathcal{M}_{t}$ exactly when $t\geq t_{m}.$ Deployment changes occur only at discrete intervals, a schedule determined by the algorithm. The algorithm partitions the time horizon $T$ into $K$ equal-length stages, where $K$ is a tunable hyperparameter of the algorithm. Each stage thus consists of $T/K$ rounds (assuming $T$ is divisible by $K$). The start of stage $k$ is $\tau_{k}=(k-1)T/K+1,k=1,\dots,K.$ Let $M_{\max}$ be the concurrency cap, i.e., the maximum number of models that can be deployed simultaneously. At each update point $\tau_{k}$, the algorithm $\mathcal{A}$ selects a deployed set $\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$ such that $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$, which then remains fixed for all $t\in[\tau_{k},\tau_{k+1}).$ Queries arriving during that stage must be routed to models in the active set $\mathcal{D}_{k}(\mathcal{A})$. Thus $\mathcal{M}_{t}$ is the available pool at time $t$, while $\mathcal{D}_{k}(\mathcal{A})$ is the active subset that can actually serve queries during stage $k$.
@@
Operational Performance and Cost. An LLM’s per-prompt quality varies with the input. However, over a long time horizon, it can be reasonably modeled as a random variable centered around a stable mean (Ding et al., 2024). Formally, each model $m$ has an unknown performance distribution $\nu_{m}(\cdot)$ supported on $[0,1]$. When $m$ is selected at time $t$, the observed score $r_{t}\in[0,1]$ is drawn from $\nu_{m}(\cdot)$ with mean $\mu_{m}\;=\;\mathbb{E}_{x\sim\nu_{m}(\cdot)}[x].$ Invoking model $m$ on a query also incurs an operational cost $c_{m_{t}}$ that combines: (i) Input Cost: $c_{m_{t}}^{(\mathrm{in})}=\bigl(\text{\# tokens of input at time }t\bigr)\times p_{\mathrm{in}},$ where $p_{\mathrm{in}}$ is the per-token input price; (ii) Output Cost: $c_{m_{t}}^{(\mathrm{out})}=\bigl(\text{\#tokens of response by }m\text{ for query }t\bigr)\times p_{\mathrm{out}},$ with output length drawn from a model-specific distribution $\xi_{m}(\cdot)$ and unit price $p_{\mathrm{out}}$.
@@
Thus the total cost is $c_{m_{t}}=c_{m_{t}}^{(\mathrm{in})}+c_{m_{t}}^{(\mathrm{out})}.$ Because the output token count $c_{m_{t}}^{(\mathrm{out})}$ depends on the specific model and query and is sampled from $\xi_{m}(\cdot)$, the total operational cost $c_{m_{t}}$ for a query handled by model $m_{t}$ is itself a random variable. Operational cost $c_{m_{t}}$ is inherently bounded by per-token pricing and practical limits on sequence length and generation (e.g., context-window and token caps). Hence we assume $c_{m_{t}}\in[c_{1},c_{2}]$ for known constants $0<c_{1}\leq c_{2}<\infty$.
@@
Constraints: Budget and Throughput. When a query arrives at time $t$, with stage index $k$ such that $\tau_{k}\leq t<\tau_{k+1}$, the algorithm $\mathcal{A}$ selects a model $m_{t}\in\mathcal{D}_{k}(\mathcal{A})$. It then observes the reward $r_{t}\sim\nu_{m_{t}}(\cdot)$ and incurs cost $c_{m_{t}}$. The goal is to maximize average reward subject to two main constraints:
(1) Budget constraint: $\mathbb{E}\bigl[\frac{1}{T}\sum_{t=1}^{T}c_{m_{t}}\bigr]\leq b+o(1).$ We use an average cost constraint instead of a hard budget constraint because for long-running systems, it provides a degree of flexibility. From a theoretical analysis perspective, since the estimation errors for both performance and cost are governed by the same concentration inequalities, the expected budget violation is upper-bounded by the order of the performance regret. This makes the average cost constraint asymptotically equivalent to a hard constraint for a near-optimal algorithm.
@@
(2) Per-model throughput limit: For each deployed LLM model $m$, we specify a throughput limit $\alpha_{m}$. Let $p_{t}(m)$ be the probability that the routing policy assigns the query at time $t$ to model $m$. While $t$ lies in stage $k$, the policy must satisfy $p_{t}(m)\leq\alpha_{m},\forall m\in\mathcal{D}_{k}(\mathcal{A}).$ This constraint caps the instantaneous load share each model may receive. For a deterministic decision $m_{t}$, where $p_{t}(m_{t})=1$, the requirement reduces to $\alpha_{m_{t}}\geq 1$. Such limits reflect real-world restrictions like API rate limits (RPM/TPM), bandwidth, or licensing, preventing any single model from being overwhelmed. Hence the deployed set’s aggregate throughput must be sufficient to serve every arrival, a condition we formalize next.
@@
The constraints are feasible. At every time $t$, there exists a subset $\mathcal{S}\subseteq\mathcal{M}_{t}$ with $|\mathcal{S}|\leq M_{\max}$ such that $\sum_{m\in\mathcal{S}}\alpha_{m}\geq 1$. The budget $b$ is also large enough to admit a non-trivial routing policy. When required, we assume Slater’s condition holds, guaranteeing strong duality for the associated optimization problems.
Performance Maximization and Regret. The goal is to maximize the expected cumulative performance, $\mathbb{E}[\sum_{t=1}^{T}\mu_{m_{t}}]$ subject to: (i) Deployment Choice: At each update $\tau_{k}$, select a deployed set $\mathcal{D}_{k}(\mathcal{A})\subseteq\mathcal{M}_{\tau_{k}}$ with $|\mathcal{D}_{k}(\mathcal{A})|\leq M_{\max}$; (ii) Model Selection: For $t\in[\tau_{k},\tau_{k+1})$, choose $m_{t}\in\mathcal{D}_{k}(\mathcal{A})$ using probabilities $p_{t}(m)$ that sum to $1$; (iii) Throughput Constraint: Ensure $p_{t}(m)\leq\alpha_{m}$ for every deployed model $m$; and (iv) Budget Constraint: Maintain $\mathbb{E}\bigl[\frac{1}{T}\sum_{t=1}^{T}c_{m_{t}}\bigr]\leq b+o(1)$.
@@
$V(b,\mathcal{S})=\max_{p\in\Delta(\mathcal{S})}\Big\{$ $\sum_{m\in\mathcal{S}}\mu_{m}p(m)\;\Big|\;\sum_{m\in\mathcal{S}}\mathbb{E}[c_{m}]p(m)\leq b,\sum_{m\in\mathcal{S}}p(m)=1,$ (1)
@@
Here, $\Delta(\mathcal{S})$ is the set of probability distributions over $\mathcal{S}$; $p(m)$ is the probability of selecting model $m$; and $\mathbb{E}[c_{m}]$ is its expected cost. The support constraint $|\text{supp}(p)|\leq M_{\max}$ limits the number of models with positive probability to $M_{\max}$, capturing the combinatorial selection of the best $M_{\max}$ models to use for routing from the entire available pool $\mathcal{M}_{\tau_{k}}$. If no feasible distribution exists or if $(\mathcal{S}=\varnothing)$, we set $V(b,\mathcal{S})=0$.
@@
3 StageRoute: Stage-Based LLM Deployment and Routing